OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Enter your answer as an ordered pair (x,y). 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (anonymous):

(3x+4y=-2)3 (2x-3y=10)4

OpenStudy (anonymous):

9x+12y=-6 8x-12y=40

OpenStudy (anonymous):

17x=34

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

(3x+4y=-2)-2 (2x-3y=10)3

OpenStudy (anonymous):

-6x-8y=4 6x-9y=30

OpenStudy (anonymous):

-17y=34

OpenStudy (anonymous):

y=-2

OpenStudy (anonymous):

(2,-2)